the bus headed north at noon at 50 mph. at 2pm, the train headed north from the same station at 70 mph. what time was it when the train got within 40 miles of the bus.
Answer by Nate(3500)

D = R*T
Bus's distance = 50t
Train's distance = 70t
The bus already traveled two hours or 100 miles.
Current Bus's distance = 50t + 100
Current Train's distance = 70t
Current Bus's distance - Current Train's distance = 40
50t + 100 - 70t = 40
100 - 20t = 40
60 = 20t
3 = t